Suppose that $3659 is deposited in a savings account which earns 6.5% simple interest.

-What is the amount due after 5 years?
-What is the amount due after 2 years if the interest is compounded?
-What is the equivalent six month interest rate?
Answer #1

Future value = Preset value ( 1 + r ) ^ n

1. Amount due after five years = 3659 (1+0.065)^5 = 5013.15

2. Amount due after 2 years if interest is compounded (assumed semi annual) = 3659 (1+0.0325)^4 = 4158.37

n = 2 x 2 = 4, r = 6.5/2 = 3.25

3. Equivalent six month interest rate = ( 1 + r/n )^n - 1 = (1+ 0.065/2)^2 -1 = 0.0661 r 6.61%

n = no of compounding r = interest rate